# The Cosmological Evidence for Non-Minimal Coupling

### v1

@@ Abstract
The recent observational evidence of deviations from the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) model points towards the presence of evolving dark energy. The simplest possibility consists of a cosmological scalar field $\varphi$, dubbed quintessence, driving the accelerated expansion. We assess the evidence for the existence of such a scalar field. We find that, if the accelerated expansion is driven by quintessence, the data favour a potential energy $V(\varphi)$ that is concave, i.e., $m^{2}=d^{2}V/d\varphi^{2}<0$. Furthermore, and more significantly, the data strongly favour a scalar field that is non-minimally coupled to gravity (Bayes factor $\log(B)=7.34\pm 0.60$), leading to time variations in the gravitational constant on cosmological scales, and the existence of fifth forces on smaller scales. The fact that we do not observe such fifth forces implies that either new physics must come into play on non-cosmological scales or that quintessence is an unlikely explanation for the observed cosmic acceleration.
@@
The simplest form of thawing dark energy – quintessence as a minimally coupled scalar field – is only marginally favoured over a cosmological constant and is not statistically favoured over the parametric CPL model of an evolving equation of state. In other words, while the data do seem to prefer a dark energy that evolves in time, the type of evolution being uncovered is not well-described by standard quintessence [12, 13]. Thus, if one is to assume that a scalar field is the source of late time acceleration, one needs to look more broadly. Effective field theory supplies us with a framework to do so.
@@
What we mean by “microphysical theory” is the action for the fundamental field that constitutes dark energy. A common assumption is that dark energy is in the form of a scalar field. Such a scalar field can generally be described by an action [25]:
@@
where $g_{\alpha\beta}$ is the metric, $R$ the Ricci scalar, $\varphi$ the scalar field, $X=\partial_{\mu}\varphi\partial^{\mu}\varphi$, and ${\cal L}_{M}$ is the action for matter fields $\psi_{M}$. $F$, $G$, $J$ and $V$ are arbitrary functions of $\varphi$.
@@
For the case of $F\neq 1$ [38, 43, 13, 44, 45, 46, 47, 14] we can consider a general subset of models by expanding $F$ and $V$,
@@
with $G(\varphi)=1$ and $J(\varphi)=0$, where throughout this paper, $\xi$ is dimensionless, $V_{0}$ is in units of $M_{\rm Pl}^{2}M^{2}_{H}$ (where $M_{H}=H_{0}/h$), $\beta$ is in units of $M_{\rm Pl}M^{2}_{H}$, and $m^{2}$ is in units of $M^{2}_{H}$. In this way, albeit keeping ourselves ignorant about the specific quintessential action, we can at least try to answer two fundamental questions about the scalar field: “what is the mass of the scalar field?” and “is it non-minimally coupled to Gravity”? [^1] We have used the freedom to shift the scalar field to remove the linear term in the non-minimal coupling.
Figure 1: Projected theory priors into the $(w_{0},w_{a})$ plane for both minimally and non-minimally coupled dark energy by fitting $(w_{0},w_{a})$ to the theory prediction of the observables used in the data constraints, accounting for their observational errors [12, 14, 48, 49].
Although restricting $F$ and $V$ to quadratic functions may seem limiting, observables are sensitive to only a narrow range of the scalar field’s evolution, making such an expansion sufficient. This approach captures a broad class of non-minimally coupled scalar field models, serving as a model-agnostic parametrization valid for any thawing scalar field with a non-derivative non-minimal coupling. For instance, it includes models with exponential and hilltop potentials studied in [38, 43, 13], as they admit a Taylor expansion like Eq. (2).
@@
We use hi_class [55, 56, 57] to compute the theory predictions, starting the field at $\varphi_{\rm ini}=0$ and ${\dot{\varphi}}_{\rm ini}=10^{-3}$ to ensure it starts rolling222 Current data only probe a small part of the potential $V(\phi)$ and are not sensitive to its full shape. In fact, it is the shape of the potential at $|\varphi|>|\varphi_{\rm ini}|$ that determines the field dynamics. As a consequence, there is little loss of generality by setting $\varphi_{\rm ini}=0$.. Finally, we tune $V_{0}$ to ensure that the Friedmann equations close (i.e., ensure that the sum of all fractional densities $\sum_{i}\Omega_{i}=1$). We adopt uniform priors on all cosmological and dark energy parameters, varying the standard cosmological parameters $\{\Omega_{b}h^{2},\Omega_{m},H_{0},\ln 10^{10}A_{s},n_{s},\tau\}$ while keeping the neutrino mass fixed to $\sum m_{\nu}=0.06\,\mathrm{eV}$. For dark energy, we consider three models: (i) Non-minimally coupled quintessence, with $\xi\in[0,3.5]$, $\beta\in[0,10]$, and $m^{2}\in[-10,10]$; (ii) Minimally coupled quintessence, corresponding to the model studied in [12, 23] where $V(\varphi)=V_{0}+\frac{1}{2}m^{2}\varphi^{2}$, with $V_{0}\in[-0.5,1.5]$ and $m^{2}\in[-100,5]$; and (iii) Parametrized dark energy, using the CPL form, with $w_{0}\in[-2,0]$ and $w_{a}\in[-2.5,1.5]$. We quantify their fit to the data by reporting the best fit $\chi^{2}=-2\log\mathcal{L}$ and its difference with respect to $\Lambda$CDM, $\Delta\chi^{2}_{X\Lambda}=\chi^{2}_{X}-\chi^{2}_{\Lambda}$. In order to account for the extra degrees of freedom, we use the Bayes factor $\log B_{X\Lambda}=\log\mathcal{Z}_{X}-\log\mathcal{Z}_{\Lambda}$ to assess the supporting evidence in favour of a given model over $\Lambda$CDM. Note that on the Jeffrey’s scale $\log B_{X\Lambda}>5$ indicates strong evidence for model $X$ relative to $\Lambda$CDM [58, 59]. The results can be found in Table 1. We consistently find that non-minimally coupled dark energy outperforms other models, with CPL coming in second, and minimally coupled dark energy coming last.
@@
$\log B_{X\Lambda}$ Min. coupled $\varphi$ $-1.96\pm 0.57$ $1.53\pm 0.57$ $3.46\pm 0.57$
$w_{0}w_{a}$ $-1.57\pm 0.60$ $4.02\pm 0.59$ $5.55\pm 0.60$
Non-min. coupled $\varphi$ $3.14\pm 0.58$ $6.45\pm 0.59$ $7.34\pm 0.60$
$\Delta\chi^{2}_{X\Lambda}$ Min. coupled $\varphi$ $-2.89$ $-7.35$ $-14.1$
$w_{0}w_{a}$ $-7.96$ $-14.0$ $-18.1$
Non-min. coupled $\varphi$ $-14.1$ $-19.4$ $-23.6$
@@
Table 1: Bayesian evidence and $\Delta\chi^{2}$ values for dynamical dark energy models compared to $\Lambda$CDM from the combination of BAO, CMB, and different SNe samples.
@@
For the CPL model, the $\Delta\chi^{2}_{w_{0}w_{a}\Lambda}$ and $\log B_{w_{0}w_{a}\Lambda}$ values are consistent with other works in the literature that have used similar datasets (e.g., [16, 15, 14, 33, 12]): in particular, $w_{0}w_{a}$ significantly improves the fit to all datasets, but is evidentially disfavoured with respect to $\Lambda$CDM when the Pantheon+ data are used, while it is evidentially favoured with Union3 and DES-Y5. For minimally coupled dark energy, while it can improve the fit to the data, it is the worst performing dynamical dark energy model, both with respect to its ability to fit the data and its Bayes factor. While it is evidentially disfavoured with respect to the Pantheon+ data (in agreement with [12, 33]); we begin to see some non-trivial evidence for it emerge with the Union3 and DES-Y5 data. Given that these datasets significantly strengthen the preference for dynamical dark energy, we attribute this improvement in evidence to the fact that this dark energy model allows for more rapid evolution at low $z$ than many of the standard, minimally coupled, quintessence models (when $m^{2}<0$) [12, 23, 28, 60]. Finally, the non-minimally coupled dark energy model provides the best fit to the data across the board. At first glance, this may not be surprising considering that it has more parameters than any of the other dark energy models considered, but the Bayesian evidence results indicate that the additional parameters justify their inclusion. The non-minimally coupled model is the only model that is evidentially favoured over $\Lambda$CDM with respect to all dataset combinations.
@@
Given that the BAO+CMB+DES-Y5 gives the strongest evidence for dynamical dark energy, we will dive a little further into this data combination. Fig. 2 depicts both the DESI DR2 BAO data for the angle-averaged distance $D_{V}$ and the DES-Y5 SNe data [^3] For visualization, the SNe data has been binned in redshift and we have calculated the weighted average distance moduli and errors using the covariance matrix to include both statistical and systematic errors. See [2, VII.B] for a similar plot. for the distance modulus $\mu$ against the best fits for all dynamical dark energy models, all normalized to the best fit $\Lambda$CDM model. Clearly, all dynamical models offer a significant improvement in fitting these distance measurements. In fact, they all perform similarly well in terms of their ability to fit the BAO and SNe data simultaneously, with $\Delta\chi^{2}_{X\Lambda}\simeq-15$ for every dynamical dark energy $X$. Where the CPL and non-minimally coupled dark energy separate themselves from minimally coupled dark energy is in their ability to simultaneously improve the fit to both the lower $z$ data and the CMB data over $\Lambda$CDM, which makes up the rest of the $\Delta\chi^{2}$ difference between these models and $\Lambda$CDM ($\Delta\chi^{2}_{X\Lambda}\simeq-3$ for CPL and $\simeq-7$ for the non-minimally coupled model).
@@
Figure 2: Observable predictions from the BAO+CMB+DES-Y5 best-fit dark energy models compared with the angle-averaged distance $D_{\rm V}$ and the distance modulus $\mu$, normalized to the $\Lambda$CDM best fit.
@@
Fig. 3 shows the constraints of the non-minimally coupled scalar field parameters. We find $\xi=2.30^{+0.71}_{-0.38}$, $\beta=2.81^{+0.50}_{-0.31}$, $m^{2}=-2.0\pm 1.3$, and $V_{0}=0.651^{+0.083}_{-0.12}$. Note that both minimally and non-minimally coupled dark energy models strongly favour $m^{2}=d^{2}V/d\varphi^{2}<0$, as concave potentials allow for more rapid variation in the scalar field [23, 12, 28, 60]. Such potentials are unbounded from below, absent of other corrections; however, this is not relevant for cosmological constraints as we find that the field excursion is $\Delta\varphi/M_{\rm Pl}\simeq{\cal O}(0.1)$ and never enters a pathological regime during the current era. Finally, the left hand panel of Fig. 4 shows the constraints on $w(z)$. As anticipated, the posterior distribution of $w(z)$ is consistently below $-1$ by more than $2\sigma$ in the matter dominated era, crossing the phantom divide at about $z\sim 0.5$ and reaching $w(z=0)=-0.841^{+0.029}_{-0.028}$ today, $5.7\sigma$ away from $w=-1$. This quick transition over this range of redshifts is what allows this model to significantly improve the fit to both the early (CMB) and late time (BAO and SNe) data, and resembles the evolution favoured by the parametric dark energy models (e.g., [16, Fig. 4]). Furthermore, we also see in the central panel that the non-minimal coupling has a significant impact on the strength of gravity, parametrized by $\mu(z)$, with $\mu(z=0)=1.77^{+0.23}_{-0.18}$, $4.3\sigma$ away from the General Relativity (GR) with $\mu=1$. These cosmological results have profound implications for the nature of gravity.
@@
Ancillary Gravitational Consequences—The non-minimal coupling of the scalar field to the metric brings with it a host of undesirable consequences and it would be remiss of us not to consider them. On the largest scales, we have that the effective gravitational constant, $G_{\rm eff}$ (as opposed to the bare gravitational constant which appears in the action, $8\pi G_{0}=M^{-2}_{\rm Pl}$) is time varying, ${\dot{G}_{\rm eff}}/G_{\rm eff}=2\xi\phi{\dot{\phi}}/(1-\xi\phi^{2})$. We find that, for parameters that can best fit the cosmological data, ${\dot{G}_{\rm eff}}/G_{\rm eff}\simeq 10^{-11}/{\rm Gyr}$. Constraints on this quantity (which are independent of any assumptions about the cosmological model) should not be confused with the constraints that come from the Solar System or compact binaries on the Newton-Poisson $G_{\rm N}$ [68, 69, 70] (although they are around the same order of magnitude [71]).
@@
Fifth forces can be suppressed on small scales, as seen from the action in Eq. (1). Adding the EFT-allowed term $\frac{1}{\Lambda^{4}}\int d^{4}x\sqrt{-g}X^{2}$ (see [43]) enables gravitational screening that protects small scales from large deviations [78, 79, 80, 81]. This leads to Vainshtein screening [82], distinct from Galileon models, due to the explicit non-minimal coupling to the Einstein-Hilbert term. The fifth force becomes sensitive to the local scalar field and, in the screened regime, takes the form $F_{5}/F_{\rm N}\propto(r/r_{\rm V})^{4/3}$ [83], where $r_{V}=[9M/(16\pi\Lambda^{2}M_{\rm Pl})]^{1/2}$, and $F_{\rm N}\propto r^{-2}$. For comparison, cubic Galileons yield $F_{5}/F_{\rm N}\propto(r/r_{\rm V})^{3/2}$ [84]. With $\Lambda\lesssim 10^{-2}\,{\rm eV}$, Solar System screening ($r_{\rm V}\gtrsim 10^{-3}\,{\rm pc}$ for $M\sim M_{\odot}$) is achievable without altering cosmic expansion. Another alternative is to break the universality of the non-minimal coupling and assume that the dark energy, in the Einstein frame, is only coupled to the dark matter. Such a scenario corresponds to an interacting dark matter/dark energy model [85, 86, 87, 88, 89, 90] and might avoid some of the constraints on fifth forces. However, as with screening, it would be another (more contrived) version of non-minimal coupling.
The only way to avoid non-minimal coupling may be through a radical reformulation of the kinetic structure of the scalar field. A purely shift-symmetric theory, replacing the kinetic term by a more complicated function, $X\rightarrow K(X,\Box\varphi)$, leads to $w<-1$ throughout and the resulting $(w_{0},w_{a})$ does not overlap with the current constraints [49]. Breaking the shift symmetry by, for example, adding a potential, may make such theories more consistent with the data, but will in turn lead to other problems, such as superluminal motion and instabilities. Such models remain to be explored.
@@
Such a non-minimal coupling will naturally arise in the effective action of a theory coupled to gravity [91, 92, 93, 94, 25, 95, 96, 97]. Furthermore, theories with extra dimensions may lead to non-minimal couplings, with one caveat: given the substantial deviation of $w(a)$ from $-1$, we inevitably have substantial field excursions. In this case we find that $\Delta\varphi/M_{\rm Pl}\sim{\cal O}(0.1)$ which is manageable, yet far too large to be accommodated in some of the better motivated theories of dark energy. Furthermore, the screening scale $\Lambda$ will be substantially larger than $(H_{0}M_{\rm Pl})^{1/2}$ and thus may lead to strong coupling problems.
@@
The quintessence hypothesis clearly leads us down a path of ever increasing complexity, and the consequences are sufficiently jarring, that it behooves us to step back and reconsider our assumptions. A key assumption is that the dark energy is in the form of a scalar field. One might want to consider the handful of proposals which do not reduce to a scalar field on large scales (or an effective description in terms of a scalar-tensor theory) [98]. Another possibility is that the details of the accelerated expansion we are inferring is a result of our mismodelling the cosmological space-time and, as a result, observables may be non-trivially affected by the inhomogeneous nature of the cosmos [99, 100, 101, 102, 103]. We should also revisit the assumptions that have gone into the analysis of the data – the presence or absence of a cosmic dipole, the role of the low redshift supernovae, the discrepancy between different supernovae and between different BAO samples, etc [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126]. We will watch with interest as the various analyses are independently scrutinized, and await the expected increase in the quality and quantity of the data from the new generation of surveys.

### v3

# Assessing cosmological evidence for non-minimal coupling

@@ Abstract
The recent observational evidence of deviations from the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) model points towards the presence of evolving dark energy. The simplest possibility consists of a cosmological scalar field $\varphi$, dubbed quintessence, driving the accelerated expansion. We assess the evidence for the existence of such a scalar field. We find that, if the accelerated expansion is driven by quintessence, the data favour a potential energy $V(\varphi)$ that is concave, i.e., $m^{2}=d^{2}V/d\varphi^{2}<0$. Furthermore, and more significantly, the data strongly favour a scalar field that is non-minimally coupled to gravity (Bayes factor $\log(B)=7.34\pm 0.6$), leading to time variations in the gravitational constant on cosmological scales, and the existence of fifth forces on smaller scales. The fact that we do not observe such fifth forces implies that either new physics must come into play on non-cosmological scales or that quintessence is an unlikely explanation for the observed cosmic acceleration.
@@
The simplest form of thawing dark energy – quintessence as a minimally coupled scalar field – is only marginally favoured over a cosmological constant and is not statistically favoured over parametric models of an evolving equation of state. In other words, while the data do seem to prefer a dark energy that evolves in time, the type of evolution being uncovered is not well-described by standard quintessence [12, 13]. Thus, if one is to assume that a scalar field is the source of late time acceleration, one needs to look more broadly.
@@
What we mean by “microphysical theory” is the action for the fundamental field that constitutes dark energy. A common assumption is that dark energy is in the form of a scalar field, which can generally be described by an action [25]:
@@
where $g_{\alpha\beta}$ is the metric, $R$ the Ricci scalar, $\varphi$ the scalar field, $X=\partial_{\mu}\varphi\partial^{\mu}\varphi$, and ${\cal L}_{M}$ is the action for matter.
@@
For the case of $F\neq 1$ [38, 43, 13, 44, 45, 46, 47, 14, 48] we can consider a general subset of models by expanding $F$ and $V$,
@@
with $G(\varphi)=1$ and $J(\varphi)=0$ (see appendix for more details on this expansion to quadratic order). Despite keeping ourselves ignorant about the specific quintessential action, we can at least try to answer two fundamental questions about the scalar field: “what is the mass of the scalar field?” and “is it non-minimally coupled to Gravity”?
Figure 1: Projected theory priors into the $(w_{0},w_{a})$ plane for both minimally and non-minimally coupled dark energy by fitting $(w_{0},w_{a})$ to the theory prediction of the observables used in the data constraints.
@@
For dark energy, we consider three models (see appendix for a discussion of parameter priors): (i) Non-minimally coupled quintessence; (ii) Minimally coupled quintessence; corresponding to the model studied in [12, 23] where $V(\varphi)=V_{0}+\frac{1}{2}m^{2}\varphi^{2}$; and (iii) the CPL parametrization. We quantify their fit to the data by reporting the best fit $\chi^{2}=-2\log\mathcal{L}$ and its difference with respect to $\Lambda$CDM, $\Delta\chi^{2}_{X\Lambda}=\chi^{2}_{X}-\chi^{2}_{\Lambda}$. In order to account for the extra degrees of freedom, we use both the $\Delta\mathrm{DIC}_{X\Lambda}$ and the Bayes factor $\log B_{X\Lambda}=\log\mathcal{Z}_{X}-\log\mathcal{Z}_{\Lambda}$ to assess the supporting evidence in favour of a given model over $\Lambda$CDM. For both $\Delta\chi^{2}$ and $\Delta\mathrm{DIC}$, negative values indicate a preference for the model, whereas for $\log B$ positive values indicate a preference for the model. Note that on the Jeffrey’s scale $\log B_{X\Lambda}>5$ indicates strong evidence for model $X$ relative to $\Lambda$CDM [56, 57]. The results can be found in Table 1, where the $\log B_{X\Lambda}$ values all have an uncertainty of 0.6$\simeq$ 0.6≃ 0.6.
@@
$\log B$
Min. $\varphi$ 1.96-1.96- 1.96 +1.53+1.53+ 1.53 +3.46+3.46+ 3.46
$w_{0}w_{a}$ 1.57-1.57- 1.57 +4.02+4.02+ 4.02 +5.55+5.55+ 5.55
Non-min. $\varphi$ +3.14+3.14+ 3.14 +6.45+6.45+ 6.45 +7.34+7.34+ 7.34
$\Delta\chi^{2}$
($\Delta\mathrm{DIC}$)
Min. $\varphi$ 2.9-2.9- 2.9 (0.32-0.32- 0.32) 7.4-7.4- 7.4 (3.41-3.41- 3.41) 14.1-14.1- 14.1 (9.96-9.96- 9.96)
$w_{0}w_{a}$ 8.0-8.0- 8.0 (4.89-4.89- 4.89) 14.0-14.0- 14.0 (12.2-12.2- 12.2) 18.2-18.2- 18.2 (16.4-16.4- 16.4)
Non-min. $\varphi$ 14.1-14.1- 14.1 (11.5-11.5- 11.5) 19.4-19.4- 19.4 (16.4-16.4- 16.4) 23.6-23.6- 23.6 (21.5-21.5- 21.5)
@@
Table 1: Statistical comparison for dynamical dark energy models relative to $\Lambda$CDM from the combination of BAO, CMB, and different SNe samples.
@@
For CPL, the $\Delta\chi^{2}_{w_{0}w_{a}\Lambda}$, $\Delta\mathrm{DIC}_{w_{0}w_{a}\Lambda}$, and $\log B_{w_{0}w_{a}\Lambda}$ values are consistent with other works in the literature that have used similar datasets (e.g., [16, 15, 14, 33, 12]): in particular, $w_{0}w_{a}$ significantly improves the fit to all datasets, but is evidentially disfavoured with respect to $\Lambda$CDM when the Pantheon+ data are used, while it is evidentially favoured with Union3 and DES-Y5. For minimally coupled dark energy, while it can improve the fit to the data, it is the worst performing dynamical dark energy model. While it is evidentially disfavoured with respect to the Pantheon+ data (in agreement with [12, 33]); we begin to see some moderate evidence for it emerge with the Union3 and DES-Y5 data. Given that these datasets significantly strengthen the preference for dynamical dark energy, we attribute this improvement in evidence to the fact that this dark energy model allows for more rapid evolution at low $z$ than standard minimally coupled quintessence models (when $m^{2}<0$) [12, 23, 28, 58]. Finally, the non-minimally coupled dark energy model provides the best fit to the data across the board. At first glance, this may not be surprising considering that it has more parameters than any of the other dark energy models considered, but the Bayesian evidence and DIC results indicate that the additional parameters justify their inclusion, with the model being strongly favoured by its Bayes factor and only receiving a modest DIC penalty. The non-minimally coupled model is the only model that is evidentially favoured over $\Lambda$CDM with respect to all dataset combinations.
@@
Given that the BAO+CMB+DES-Y5 gives the strongest evidence for dynamical dark energy, we will dive a little further into this data combination. Fig. 2 depicts both the DESI DR2 BAO data for the angle-averaged distance $D_{V}$ and the DES-Y5 SNe data for the distance modulus $\mu$ against the best fits for all dynamical dark energy models, all normalized to the best fit $\Lambda$CDM model. For visualization, the SNe data has been binned in redshift and we have calculated the weighted average distance moduli and errors using the covariance matrix to include both statistical and systematic errors. See [2, VII.B] for a similar plot. Clearly, all dynamical models offer a significant improvement in fitting these distance measurements. In fact, they all perform similarly well in terms of their ability to fit the BAO and SNe data simultaneously, with $\Delta\chi^{2}_{X\Lambda}\simeq-15$ for every dynamical dark energy model. Where the CPL and non-minimally coupled dark energy separate themselves from minimally coupled dark energy is in their ability to simultaneously improve the fit to both the lower $z$ data and the CMB data over $\Lambda$CDM, which makes up the rest of the $\Delta\chi^{2}$ difference between these models and $\Lambda$CDM ($\Delta\chi^{2}_{X\Lambda}\simeq-3$ for CPL and 7$\simeq$-7≃ - 7 for the non-minimally coupled model).
@@
Figure 2: Observables from the best-fit dark energy models compared with the angle-averaged distance $D_{\rm V}$ and the distance modulus $\mu$, normalized to $\Lambda$CDM.
@@
Fig. 3 shows the constraints of the non-minimally coupled scalar field parameters. We find $\xi=2.31^{+0.75}_{-0.34}$, $\beta=2.85^{+0.53}_{-0.29}$, $m^{2}=-1.8\pm 1.2$, and $V_{0}=0.654^{+0.086}_{-0.13}$. Note that both minimally and non-minimally coupled dark energy models strongly favour $m^{2}=d^{2}V/d\varphi^{2}<0$, as concave potentials allow for more rapid variation in the scalar field [23, 12, 28, 58]. Such potentials are unbounded from below, absent of other corrections; however, this is not relevant for cosmological constraints as we find that the field excursion is $\Delta\varphi/M_{\rm Pl}\simeq{\cal O}(0.1)$ and never enters a pathological regime. Finally, the left hand panel of Fig. 4 shows the constraints on $w(z)$. As anticipated, the posterior distribution of $w(z)$ is consistently below 1-1- 1 by more than $2\sigma$ in the matter dominated era, crossing the phantom divide at about $z\sim 0.5$ and reaching $w(z=0)=-0.841^{+0.029}_{-0.028}$ today, $5.7\sigma$ away from $w=-1$. This quick transition over this range of redshifts is what allows this model to significantly improve the fit to both the early (CMB) and late time (BAO and SNe) data simultaneously, and resembles the evolution favoured by the parametric dark energy models (e.g., [16, Fig. 4]). Furthermore, we also see in the central panel that the non-minimal coupling has a significant impact on the strength of gravity, parametrized by $\mu(z)$, with $\mu(z=0)=1.77^{+0.23}_{-0.18}$, $4.3\sigma$ away from the General Relativity (GR) with $\mu=1$. These cosmological results have profound implications for the nature of gravity.
@@
Ancillary Gravitational Consequences—The non-minimal coupling of the scalar field to the metric brings with it a host of undesirable consequences and it would be remiss of us not to consider them. On the largest scales, we have that the effective gravitational constant, $G_{\rm eff}$ (as opposed to the bare gravitational constant which appears in the action, $8\pi G_{0}=M^{-2}_{\rm Pl}$) is time varying, ${\dot{G}_{\rm eff}}/G_{\rm eff}=2\xi\varphi{\dot{\varphi}}/(1-\xi\varphi^{2})$. We find that, for parameters that can best fit the cosmological data, ${\dot{G}_{\rm eff}}/G_{\rm eff}\simeq 10^{-11}/{\rm yr}$. Constraints on this quantity (which are independent of any assumptions about the cosmological model) should not be confused with the constraints that come from the Solar System or compact binaries on the Newton-Poisson $G_{\rm N}$ [66, 67, 68] (although they are around the same order of magnitude [69]).
@@
Fifth forces can be suppressed on small scales, as seen from the action in Eq. (1). Adding the EFT-allowed term $\frac{1}{\Lambda^{4}}\int d^{4}x\sqrt{-g}X^{2}$ enables gravitational screening that protects small scales from large deviations [76, 77, 78, 79, 43, 80]. This leads to Vainshtein screening [81], distinct from Galileon models, due to the explicit non-minimal coupling to the Einstein-Hilbert term. The fifth force becomes sensitive to the local scalar field and, in the screened regime, takes the form $F_{5}/F_{\rm N}\propto(r/r_{\rm V})^{4/3}$ [82], where $r_{V}=[9M/(16\pi\Lambda^{2}M_{\rm Pl})]^{1/2}$, and $F_{\rm N}\propto r^{-2}$. For comparison, cubic Galileons yield $F_{5}/F_{\rm N}\propto(r/r_{\rm V})^{3/2}$ [83]. With $\Lambda\lesssim 10^{-2}\,{\rm eV}$, Solar System screening ($r_{\rm V}\gtrsim 10^{-3}\,{\rm pc}$ for $M\sim M_{\odot}$) is achievable without altering cosmic expansion. Another alternative is to break the universality of the non-minimal coupling and assume that the dark energy, in the Einstein frame, is only coupled to the dark matter. Such a scenario corresponds to an interacting dark matter/dark energy model [84, 85, 86, 87, 88, 89, 90] and might avoid some of the constraints on fifth forces. However, as with screening, it would be another (more contrived) version of non-minimal coupling.
The only way to avoid non-minimal coupling may be through a radical reformulation of the kinetic structure of the scalar field. A purely shift-symmetric theory, replacing the kinetic term by a more complicated function, $X\rightarrow K(X,\Box\varphi)$, leads to $w<-1$ throughout and the resulting $(w_{0},w_{a})$ does not overlap with the current constraints [50]. Breaking the shift symmetry by, for example, adding a potential, may make such theories more consistent with the data, but will in turn lead to other problems, such as superluminal motion and instabilities. We will explore these models in future work.
@@
Such a non-minimal coupling will naturally arise in the effective action of a theory coupled to gravity [91, 92, 93, 94, 25, 95, 96, 97]. Furthermore, theories with extra dimensions may lead to non-minimal couplings, with one caveat: given the substantial deviation of $w(a)$ from 1-1- 1, we inevitably have substantial field excursions. In this case we find that $\Delta\varphi/M_{\rm Pl}\sim{\cal O}(0.1)$ which is manageable, yet far too large to be accommodated in some of the better motivated theories of dark energy. Furthermore, the screening scale $\Lambda$ will be substantially larger than $(H_{0}M_{\rm Pl})^{1/2}$ and thus may lead to strong coupling problems.
@@
The quintessence hypothesis clearly leads us down a path of ever increasing complexity, and the consequences are sufficiently jarring, that it behooves us to step back and reconsider our assumptions. A key assumption is that the dark energy is in the form of a scalar field. One might want to consider the handful of proposals which do not reduce to a scalar field on large scales (or an effective description in terms of a scalar-tensor theory) [98]. Another possibility is that the details of the accelerated expansion we are inferring is a result of our mismodelling the cosmological space-time and, as a result, observables may be non-trivially affected by the inhomogeneous nature of the cosmos [99, 100, 101, 102, 103]. We should also revisit the assumptions that have gone into the analysis of the data – the presence or absence of a cosmic dipole, the role of the low redshift supernovae, the discrepancy between different supernovae and between different BAO samples, etc [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131]. We will watch with interest as the various analyses are independently scrutinized, and await the expected increase in the quality and quantity of the data from the new generation of surveys.